A man bought 2 pounds of coffee and 1 pound of butter for a total of $18.75. A month later the prices had not change (this makes it a fictitious problem), and he bought 3 pounds of coffee and 2 pounds of butter for $29.50. Find the price per pound of both the coffee and the butter.
Answer by checkley71(8403)

2c+1b=18.75 multiply this equation by -2 & add them
3c+2b=29.50
-4c-2b=-37.50
--------------------------
-c=-8
c=8 price if the coffee is $8.00
2*8+b=18.75
16+b=18.75
b=18.75-16
b=2.75 price of the butter.
proof
3*8+2*2.75=29.50
24+5.50=29.50
29.50=29.50